OpenStudy (anonymous):

Two bags containing 88 and 112 jelly beans were combined. The percent of watermelon jelly beans in each bag was 25% and 50%, respectively. Anthony selects one jelly bean. What is the probability he selected a watermelon jelly bean?

OpenStudy (foolaroundmath):

Number of watermelon jelly beans in first bag = 25% of 88 = 88*25/100 = 22 Number of watermelon jelly beans in second bag = 50% of 112 = 0.5*112 = 56 Total number of watermelon jelly beans = 22+56 = 78 Total number of jelly beans = 88+112 = 200 So, probability of selecting a watermelon jelly bean = 78/200 = 0.39

OpenStudy (anonymous):

thanks!